OpenStudy (anonymous):

Simplify the following expression: 2x - 2y + 5z - 2x - y + 3z x - 3y + 8z 2x - 2y + 8z 3y + 8z -3y + 8z

OpenStudy (anonymous):

I can help you if you re-structure the question in an easier way to understand. Is the entire line one expression or are their several expressions?

OpenStudy (anonymous):

Your conclusion is correctly simplified, however your steps are a little faulty

OpenStudy (anonymous):

-2+8z

OpenStudy (anonymous):

it is because 2x - 2y + 5z - 2x - y + 3z 2x-2x=0 -2y-y=-2 5z+3z=8z -2+8z

OpenStudy (anonymous):

Simplify the following expression: 2x - 2y + 5z - 2x - y + 3z

OpenStudy (anonymous):

yeah